40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 25 15:33:55 2012, janosch, Update, LSC, ringdown measurement Rindown1.png
    Reply  Wed Jul 25 16:31:01 2012, Jenne, Update, LSC, Yarm green locking to arm - PDH box saturating 
    Reply  Wed Jul 25 22:16:36 2012, Koji, Update, LSC, ringdown measurement 
Message ID: 7030     Entry time: Wed Jul 25 16:31:01 2012     In reply to: 7029
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Yarm green locking to arm - PDH box saturating 

Quote:

... it was not possible to get the cavity locked in green. So we decided to do the first measurement with infrared locked only. 

 When we sat down to align the Yarm to the green, the green light was happy to flash in the cavity, but wouldn't lock, even after Jan had tweaked the mirrors such that we were flashing the TEM00 mode.  When we went down to the end to investigate, the Universal PDH box was saturating both negative and positive.  Turning down the gain knob all the way to zero didn't change anything, so I put it back to 52.5.  Curiously, when we unplugged the Servo OUT monitor cable (which was presumably going to the rack to be acquired), the saturation happened much less frequently.  I think (but I need to look at the PDH box schematic) that that's just a monitor, so I don't know why that had to do with anything, but it was repeatable - plug cable in, almost constant saturation....unplug cable, almost no saturation.

Also, even with the cable unplugged, the light wouldn't flash in the cavity.  When I blocked the beam going to the green REFL PD (used for the PDH signal), the light would flash.

Moral of the story - I'm confused.  I'm going to look up the PDH box schematic before going back down there to investigate.

ELOG V3.1.3-